Find the magnitude and direction of the following

2-space vectors.

a. (−3, 4)

b. (2, 7)

c. (−5, −3)


1
Expert's answer
2021-04-29T17:28:23-0400

Magnitude of the 2- space vectors are


a) For (-3,4) the answer is "\sqrt{9+16}=\sqrt{25}=5" .


b) For (2,7) the answer is "\sqrt{4+49}=\sqrt{53}"


c) For (-5,-3) the answer is "\sqrt{25+9}=\sqrt{34}"



Direction of the 2-space vectors are


a) "tan^{-1}(\frac{4}{-3})"


b) "tan^{-1}{\frac{7}{2}}"


c) "tan^{-1}{\frac{-3}{-5}}=tan^{-1}{\frac{3}{5}}"